The javascript on this page is used to preload the images for links and to swap the images used for links
go to content
layout image go to introduction
go to science datago to mission status reportgo to location / positiongo to 3d data mapsgo to moon VRMLgo to further reading
layout image
Layout imagego to home
go to projectgo to resultsGo to datavisgo to historyGo to scienceGo to educationGo to archivesgo to resources


Lunar Prospector Status Report #55

October 29, 1998 - 1:00 p.m. EST (10:00 a.m. PST)

Lunar Prospector Mission Status Report #55 The Lunar Prospector spacecraft continues to perform very well.

The transmitter continues to be cycled during occultations when the propellant tank heater is on. There was one commanding problem at the station which caused a 28 minute delay in transmission of the transmitter off/on commands on DOY 300. Because we were in a period of operations without occultations, there was no loss of data.

On DOY 300, at 1757 GMT, some of the data from the spectrometer instruments became garbled. Analysis by the science team at LANL suggested a software error in the spectrometer electronics. 35 commands were sent to reset the electronics power and reconfigure the instruments. They were all reconfigured and producing good data at 302/0034. On 302 at 1615 GMT, 2 commands were sent to tweak the gain of the NS instrument.


Current spacecraft state (0000 GMT 10/29/98):

Orbit: 3536
Downlink: 3600 bps
Spin Rate: 12.00 rpm

Spin Axis Attitude (ecliptic):
Latitude:-89.0 deg
Longitude: 166 deg

Trajectory:
Periapsis Alt: 87.1 km
Apoapsis Alt: 112.9 km
Period: 118 min

Occultations: 32 minutes
Eclipses: 45 minutes
Propellant Remaining: 24.1 kg